If for all real values of `a`, one zero of the polynomial x^2-3ax+f(a) is double of the other zero then find f(x).

Grade:9

1 Answers

Vikas TU
6 years ago
Dear Student,
let one root be m , then the other root must be 2m. 
Sum of roots = 3m = 3a 
M= a 
Product of roots = 2m^2 = 2a^2 = f(a) 
F(x)= 2x^2 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)